OpenStudy (anonymous):

What is the percent decrease modeled in the function, f(x)=0.45(0.91)^x? 55% 45% 9% 91%

OpenStudy (solomonzelman):

\(\large\color{slate}{ f(x)=0.45(0.91)^x }\) that means that you multiply times \(\large\color{slate}{ 0.91 }\) each time, correct?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

multiplying times \(\large\color{slate}{ 0.91 }\) means that you are taking \(\large\color{slate}{ 91\text{%} }\) of what you previously had. When you take \(\large\color{slate}{ 91\text{%} }\), that means that you leave out the other \(\large\color{slate}{ 9\text{%} }\)...

OpenStudy (solomonzelman):

does that make sense?

OpenStudy (anonymous):

Yes it does but why exactly does it have to be 91%?

OpenStudy (anonymous):

What about the 0.45?

OpenStudy (solomonzelman):

well, because multiplying times \(\large\color{slate}{ 0.91 }\) is same as taking \(\large\color{slate}{ 91\text{%} }\). conversion from numbers to percent \(\large\color{slate}{ \color{teal}{0.91}=(0.91\times 100)\text{%} =\color{teal}{91\text{%}} }\) \(\large\color{slate}{ \color{teal}{0.44}=(0.44\times 100)\text{%} =\color{teal}{44\text{%}} }\) \(\large\color{slate}{ \color{teal}{0.03}=(0.03\times 100)\text{%} =\color{teal}{3\text{%}} }\)

OpenStudy (solomonzelman):

what do you mean what about 0.45?

OpenStudy (anonymous):

0.45(0.91)^x so does that 0.45 mean nothing?

OpenStudy (solomonzelman):

that is just the coefficient, because, for any function: \(\large\color{slate}{ y=a(b)^x}\) when \(\large\color{slate}{ x=2}\), then \(\large\color{slate}{ y=a \times (b^2)}\) when \(\large\color{slate}{ x=3}\), then \(\large\color{slate}{ y=a \times (b^3)}\) when \(\large\color{slate}{ x=4}\), then \(\large\color{slate}{ y=a \times (b^4)}\) and on...

OpenStudy (solomonzelman):

so, \(\large\color{slate}{ 0.45(0.91)^x}\) would mean that as you add 1 to the X, you multiply another time by 0.91

OpenStudy (anonymous):

What is the percent increase modeled in the function f(x)=20(1.5)^x ? what about the increase?

OpenStudy (solomonzelman):

can you convert 1.5 into the percent for me?

OpenStudy (solomonzelman):

(multiply times 100, as I did before in 3 of my examples)

OpenStudy (anonymous):

the percentage is 150%

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

1.5 is same as 150 percent.

OpenStudy (solomonzelman):

even if I am typing don't be afraid to interrupt me, by posting anything....

OpenStudy (solomonzelman):

so \(\large\color{slate}{ 150\text{%} }\) is how much percent more than \(\large\color{slate}{ 100\text{%} }\) ?

OpenStudy (anonymous):

Alright, I am really confused and would be 50%

OpenStudy (solomonzelman):

yes, that would be 50%.

OpenStudy (solomonzelman):

that means that you have a so \(\large\color{slate}{ 50\text{%} }\) increase in your case.

OpenStudy (solomonzelman):

so \(\large\color{slate}{ 100\text{%} }\) is the entire thing. when you take so \(\large\color{slate}{ 150\text{%} }\) you are taking an additional so \(\large\color{slate}{ 50\text{%} }\) on that.

OpenStudy (anonymous):

I am starting to understand that now somethings that confuse me is the graph here it is http://prntscr.com/5ng1tr

OpenStudy (solomonzelman):

okay, if you flipped the blue graph (I mean reflected across the x axis), would it be an exponential increase, or an exponential decay?

OpenStudy (anonymous):

increase

OpenStudy (solomonzelman):

yes. And when you flip this exponential increase, you are doing it by adding a - to it, because you will have all of the same outputs (y values) as the exponential increase function, except that they all will be negative.

OpenStudy (solomonzelman):

which of the options given, is a normal exponential increase function, THAT IS REFLECTED ACROSS THE X AXIS, by adding a minus to the coefficient?

OpenStudy (solomonzelman):

lost?

OpenStudy (anonymous):

Yes kinda.

OpenStudy (solomonzelman):

okay, lets consider this:  the left side picture describes the right side picture describes \(\large\color{slate}{ y=a(b)^x }\) (when b>1) \(\large\color{slate}{ y=-a(b)^x }\) (when b>1)